OpenStudy (anonymous):

if the sides of a square are lengthend by 8cm, the area becomes 169cm^2. find the lenght of a side of the original square.

OpenStudy (radar):

(x+8cm)^2=169 x+8=13 x=5

OpenStudy (anonymous):

let side of the original square x the side of the new square will be 8+x

OpenStudy (anonymous):

and (x+8)^2=169 then x=5

OpenStudy (anonymous):

just like what radar said

OpenStudy (radar):

Yes, and the negative result is not applicable.

OpenStudy (anonymous):

oh k

OpenStudy (radar):

I did not show the negative answer but you realize that the square root of 169 is -13 as well as +13.

OpenStudy (radar):

\[\pm13\]

OpenStudy (radar):

So a negative answer would of been -21 but senseless.

OpenStudy (radar):

Something very important, which I did not show, but mukushia showed you how to set the problem up.

OpenStudy (anonymous):

ok so i got that the original length was 5cm am i right

OpenStudy (radar):

Yes indeed.